# Interactive Query Answering on Knowledge Graphs with Soft Entity Constraints

### v3

@@ 6.3 Baselines
We consider a traditional Learning-to-Rank (LTR) baseline based on LightGBM (ke2017lightgbm), a gradient boosting framework implementing the LambdaRank objective (burges_learning_2005; burges_learning_2006). For each training query, the model receives as input three scalar features for every entity $e$: the original score $a[e]$, and the average positive and negative similarities, $\delta(P^{+},e)$ and $\delta(P^{-},e)$. Graded relevance labels are assigned following standard LTR practice (burges2010ranknet): entities in $P^{+}$ receive a label of 2, entities in $P^{-}$ receive 1, and a random set of 100 negative entities is assigned 0. The model is trained to predict ranking scores that order entities according to their relative relevance by directly optimizing NDCG via the LambdaRank objective. This baseline allows us to compare our proposed linear score update based on shifts from the original score, with an approach that predicts the score using a nonlinear function of relevant features.
6.4 Experimental details
For all models, we find the best values of hyperparameters using the validation set via grid search. In all cases, we use QTO (bai2023qto) as the base QA model. More details about training and hyperparameter tuning can be found in Appendix D. We implement all of our experiments using PyTorch (paszke_pytorch_2019), and run them on a machine with one NVIDIA A100 GPU and 120GB of RAM.
@@ 7 Results
The NDCG@10 results provide a unified view of the two underlying criteria: pairwise ordering and overall ranking quality. For small preference sets, all constrained models achieve higher NDCG@10 than the Unconstrained ranking, primarily due to better pairwise accuracy. Since both PA and ranking metrics improve as the size of the preference sets increases, NDCG@10 exhibits the same upward trend. Importantly, NDCG@10 allows comparing which method best satisfies the dual objective of query answering under soft constraints. We observe that for small preference sets, NQR performs best, suggesting that its learned conditional weights help preserve ranking quality while applying fine-grained score adjustments. For larger sets, the Cosine update surpasses NQR, matching or slightly improving ranking quality while achieving higher PA. This indicates that Cosine is a suitable general-purpose update rule, whereas NQR is preferable when fewer preferences are available.
We present additional results in Appendix˜E where metrics are categorized according to each of the types of complex queries in the datasets. The results show that on average Cosine performs best at balancing similarity constraints with global ranking performance, followed by NQR, which for certain query types, preserves best global ranking performance.
@@
Figure˜4 illustrates how pairwise accuracy (PA) and mean reciprocal rank (MRR) evolve jointly for different values of $\alpha$ and $\beta$ in the Cosine update (Equation˜10), as the number of interaction steps increases. Across both datasets, increasing $\alpha$ (which implies placing more weight on the original scores), prevents significant drops in MRR, but also limits the maximum PA that can be reached. Smaller $\alpha$ values, on the other hand, allow stronger preference propagation, and the effect on MRR varies depending on the value of $\beta$. This behavior reflects an important trade-off: higher $\alpha$ yields conservative and stable adjustments, while lower $\alpha$ enables more aggressive shifts guided by similarity signals.
@@
Table˜3 shows a 2-hop query from FB15k-237, where the task is to rank awards associated with people born in New York City. Applying similarity constraints on the target variable using the Cosine update sharply improves ranking quality: literary awards in the positive set are promoted into the Top-5, while unrelated film and TV awards drop in rank. NDCG@10 increases from 0.0 to 0.60, illustrating how a small preference set can adjust the ranking towards the soft constraint that awards returned as an answer to the query should be literary rather than film or music awards.
Table 2: Average execution time per query for different reranking methods on FB15k237.
@@
We present the average runtime in milliseconds per query in the FB215k237 dataset for all methods in Table˜2, where we include the added runtime ($\Delta$) with respect to the Unconstrained baseline. We note that LightGBM results in very high overhead, by adding 13.8 ms per query, due to the use of sequential tree traversal. On the other hand, the Cosine and NQR methods add very little overhead of up to 1.4 ms per query. While the average runtime of Cosine seems higher than NQR, a paired t-test does not show that the difference is statistically significant.
@@
Initial Top-5 Top-5 after Cosine update
Golden Globe Award for Best Original Score $\blacktriangle$ Nebula Award for Best Novel
Academy Award for Best Original Song $\blacktriangle$ Hugo Award for Best Novella
Academy Award for Best Director $\blacktriangle$ Hugo Award for Best Short Story
Academy Award for Best Original Screenplay $\blacktriangle$ Hugo Award for Best Professional Artist
Nobel Prize in Literature $\blacktriangle$ Hugo Award for Best Novel
@@ Appendix C Dataset statistics
We consider 14 types of complex queries over KGs introduced in prior work (ren_beta_2020), which include conjunctions, disjunctions, and negation. These queries can be represented using query graphs, which we illustrate in Fig. 5.
@@ Appendix D Experimental details
We make use of QTO as the base query answering model bai2023qto. QTO relies on a neural link predictor, for which we use ComplEx (trouillon2016complex; lacroix2018complex_n3) trained with an auxiliary relation prediction objective (chen_relation_2021). We use an embedding dimension of 1,000, learning rate of 0.1, batch size 1,000, and a regularization weight of 0.05.
@@
We search for the best values of $\alpha_{p},\alpha_{n}$ via a grid search where each can take values in $\{0.1,0.25,0.5,0.75,0.9\}$, based on performance on the validation set. We define performance as the sum of pairwise accuracy and ranking performance.
@@
We describe the architecture of NQR in Table 5. When using the RankNet loss we select the best learning rate in $\{1\times 10^{-5},1\times 10^{-4},1\times 10^{-3}\}$. For NQR, we perform a grid search where the learning rate takes values in $\{1\times 10^{-5},1\times 10^{-4}\}$, the margin in the preference loss in $\{0.05,0.1,0.25\}$, and the KL divergence weight in $\{0.1,1.0,10\}$.
Table 5: Layer-wise description of the model architecture in NQR. $t$ indicates the number of input preferences, which is variable. The embedding dimension is 1,000, and in specific layers we concatenate the score of an entity, indicated as +1+1.
@@
Preference embedding
Self-attention $(t,1000+1)$ $(t,1001)$
LayerNorm $(t,1001)$ $(t,1001)$
Linear + ReLU (fc1) $(t,1001)$ $(t,1000)$
Mean-pooling $(t,1000)$ 10001000
Score adjustment
Linear + ReLU $(2000+1)$ 10001000
Linear + tanh 10001000 11
@@ Appendix E Additional results
We present here results over all types of complex queries we consider in our experiments. As in our main experiments, we compute the metrics for preference sets of increasing sizes, from 1 up to 10. Here, we report the metrics averaged over all 10 steps. We present results categorized by query structure (illustrated in Fig. 5) in Tables 6 and 7.

### v4

@@ 6.3 Baselines
We consider a traditional Learning-to-Rank (LTR) baseline based on LightGBM (Ke et al., 2017), a gradient boosting framework implementing the LambdaRank objective (Burges et al., 2005; 2006). For each training query, the model receives as input three scalar features for every entity $e$: the original score $a[e]$, and the average positive and negative similarities, $\delta(P^{+},e)$ and $\delta(P^{-},e)$. Graded relevance labels are assigned following standard LTR practice (Burges, 2010): entities in $P^{+}$ receive a label of 2, entities in $P^{-}$ receive 1, and a random set of 100 negative entities is assigned 0. The model is trained to predict ranking scores that order entities according to their relative relevance by directly optimizing NDCG via the LambdaRank objective. This baseline allows us to compare our proposed linear score update based on shifts from the original score, with an approach that predicts the score using a nonlinear function of relevant features. For all models, we find the best values of hyperparameters using the validation set via grid search. In all cases, we use QTO (Bai et al., 2023) as the base QA model. More details about training and hyperparameter tuning can be found in Appendix D.
@@ 7 Results
The NDCG@10 results provide a unified view of the two underlying criteria: pairwise ordering and overall ranking quality. For small preference sets, all constrained models achieve higher NDCG@10 than the Unconstrained ranking, primarily due to better pairwise accuracy. Since both PA and ranking metrics improve as the size of the preference sets increases, NDCG@10 exhibits the same upward trend. Importantly, NDCG@10 allows comparing which method best satisfies the dual objective of query answering under soft constraints. We observe that for small preference sets, NQR performs best, suggesting that its learned conditional weights help preserve ranking quality while applying fine-grained score adjustments. For larger sets, the Cosine update surpasses NQR, matching or improving ranking quality while achieving higher PA.
We present additional results in Appendix˜E where metrics are categorized according to each of the types of complex queries in the datasets. The results show that on average Cosine performs best at balancing similarity constraints with global ranking performance, followed by NQR, which for certain query types, preserves best global ranking performance. Furthermore, in we present results on manually curated preference sets, in which we confirm the trends we observe when evaluating models on preference sets derived from embeddings, supporting their use to estimate human notions of similarity.
Discussion.
Task-specific training of NQR is particularly beneficial for preference sets of size 4 or smaller, as it prevents large drops in ranking performance (MRR) while accurately capturing preferences (pairwise accuracy). This is especially relevant in settings where preference sets are costly to obtain, such as human annotation. Moreover, higher NDCG@10 for preference sets up to size 8 across both datasets suggests that NQR achieves a better balance between ranking quality and preference alignment.
These results highlight a key challenge in incorporating soft constraints into complex queries. Pairwise accuracy reflects whether preferred entities are ranked above non-preferred ones, whereas ranking metrics depend on the absolute positions of correct answers. With few preferences, the reranking signal is sufficient to separate preferred from non-preferred entities locally, improving pairwise accuracy, but not strong enough to consistently reorder the full ranking. This can temporarily promote non-answer entities, leading to lower MRR despite improved pairwise ordering. As more preferences are incorporated, the similarity signal becomes more stable and better aligned with the structure of the answer set. This reduces such spurious promotions and allows the ranking to recover, ultimately improving both preference alignment and global ranking metrics.
@@ Discussion.
Figure˜5 illustrates how pairwise accuracy (PA) and mean reciprocal rank (MRR) evolve jointly for different values of $\alpha$ and $\beta$ in the Cosine update (Equation˜8), as the number of interaction steps increases. Across both datasets, increasing $\alpha$ (which implies placing more weight on the original scores), prevents significant drops in MRR, but also limits the maximum PA that can be reached. Smaller $\alpha$ values, on the other hand, allow stronger preference propagation, and the effect on MRR varies depending on the value of $\beta$. This behavior reflects an important trade-off: higher $\alpha$ yields conservative and stable adjustments, while lower $\alpha$ enables more aggressive shifts guided by similarity signals.
@@
Table˜3 shows a 2-hop query from FB15k237, where the task is to rank awards associated with people born in New York City. Applying similarity constraints on the target variable using the Cosine update sharply improves ranking quality: literary awards in the positive set are promoted into the Top-5, while unrelated film and TV awards drop in rank. NDCG@10 increases from 0.0 to 0.60, illustrating how a small preference set can adjust the ranking towards the soft constraint that awards returned as an answer to the query should be literary rather than film or music awards.
We also observe a failure mode in which the process of reranking can cause correct answers to be demoted below incorrect entities. We show the affected entities, together with their drop in position from the original ranking.
Table 2: Average execution time per query for different reranking methods and difference ($\Delta$) with Unconstrained.
@@
This highlights the trade-off that arises when incorporating soft constraints: reranking improves alignment with the soft constraint, but can also over-propagate similarity signals and harm some correct answers, especially with limited preferences.
@@
We present the average runtime in milliseconds per query in the FB215k237 dataset for all methods in Table˜2, where we include the added runtime ($\Delta$) with respect to the Unconstrained baseline. We note that LightGBM results in very high overhead, by adding 13.8 ms per query, due to the use of sequential tree traversal.
On the other hand, the Cosine and NQR methods add very little overhead of up to 1.4 ms per query. While the average runtime of Cosine seems higher than NQR, a paired t-test does not show that the difference is statistically significant.
@@
Initial Top-5 Top-5 after update Incorrectly demoted answers
[…]
@@
Reranking with large language models. Finally, we consider the use of large language models (LLMs) when given access to textual descriptions (which we do not assume in our reranking experiments). Importantly, LLMs are limited by the high computational cost required to encode text where memory consumption increases quadratically (Vaswani et al., 2017). To quantify this, we measure the runtime of an LLM when clustering the set of answers to a query (instead of reranking all entities in the KG, as our experiments require), and we find that the average runtime per answer set is 3.5 seconds in average over 140 queries (95% CI [3.3, 3.8]). While this includes network effects, it is much larger than the slower method in our experiments which require reranking betweeen 14k to 45k entities per query. This motivates future work on using LLMs for incorporating soft constraints on queries over large KGs.
@@ Appendix B Generating preference data
Validation.
The use of embedding models as proxies of similarity is supported by prior works indicating their correlation with human judgments (Reimers and Gurevych, 2019; Grand et al., 2022; Merkx et al., 2022). In order to quantify this correlation in the context of preference sets generated via embeddings, we build a set of manually curated preference sets for queries in the FB15k237 dataset, and we then measure how well they overlap with our preference sets derived from embeddings. This process consists of three steps:
  1. 1.
    Candidate generation. Generating partitions of query answer sets can be a very challenging task, as a single query can have tens to hundreds of answers and several different ways to define a partition, resulting in very long annotation times. To make this study feasible, we rely on a large language model (LLM) to propose an initial candidate partition, which is then approved by a human. In particular, for a given query and its set of answers, we prompt an LLM (GPT 5.1) to propose a partition of the answer set into positives and negatives by relying on a vague property that is hard to formalize. The full prompt is as follows:
    “You are clustering entities by a human-interpretable shared property that is difficult to formalize. Given candidate entity descriptions, choose one coherent positive group and assign the rest to negatives. Prefer positives that share a specific latent theme, not a generic category that can be easily found in a knowledge graph. Only rely on the explicit information provided about each entity, not on background knowledge. Return only valid JSON with keys: label, rationale, positives, negatives. The positives and negatives must be lists of integer candidate indices. Positives must be non-empty, negatives must be non-empty, and together they must cover each candidate exactly once.”
  2. 2.
    Manual selection. A human (co-author in this study) is asked to confirm whether each candidate is a valid partition of the answer set. The annotation guideline is the following: “a clustering is valid if the entities in the positive set share a common property not represented by the entities in the negative set, and that is difficult to formalize with first-order logic.” We repeat this process until we collect preference sets for 10 queries of each of the 14 types we consider in our work (see Figure˜6), for a total of 140 queries. We present examples of the rationales for the selected preference sets in Table˜4.
  3. 3.
    Agreement evaluation. For each of the 140 queries, we compare the embedding-based partition of its answer set with the manually curated partition. Treating the manual annotation as ground truth, we compute accuracy as the fraction of entities whose positive/negative assignment matches between the two partitions. Across all 140 annotated queries, the mean accuracy between the curated preference sets and the embedding-based ones is 0.7904, with 95% CI [0.7677, 0.8129]. Table 5 shows the breakdown by query type. Agreement is strongest for path/intersection families such as ip (0.8778), 2i (0.8663), 1p (0.8560), and 2p (0.8498), while union and negation-heavy families remain more challenging, notably 2u-DNF (0.6539) and pni (0.6611).
The observed accuracy thus indicates a substantial match between preference sets derived from embeddings and from human notions of similarity, supporting the use of synthetic clusters as a practical proxy for training and evaluating models at the scale of thousands of queries, where fully manual annotation would be prohibitively expensive.
Mid-sized interior Western US cities (not state capitals) near mountains or high plains
Primarily known as record producers / mixing engineers rather than as front-facing recording artists
Films adapted from literary works that are not straightforward biographies or autobiographies
Primarily visual artists rather than musicians/performers
American screenwriters/playwrights primarily known for writing (not acting) for film/TV or stage
Primarily celebrated as innovative lead guitarists in major rock bands rather than as solo singer‑songwriters
Biographical dramas about real historical/political figures (film or TV)
American film/TV composers (primarily known for composing scores rather than performing)
Rock subgenres defined by blending rock with electronic/technological elements or dance music
Writers strongly associated with speculative genres (fantasy, science fiction, or horror)
Table 4: Examples of human-validated rationales for preference sets.
Query Type Accuracy 95% CI
1p 0.8560 [0.7687, 0.9410]
2p 0.8498 [0.7663, 0.9234]
3p 0.8238 [0.7448, 0.9009]
2i 0.8663 [0.7963, 0.9361]
3i 0.7175 [0.6197, 0.8068]
ip 0.8778 [0.8097, 0.9436]
pi 0.8376 [0.7638, 0.9038]
2in 0.7640 [0.7003, 0.8344]
3in 0.7588 [0.6825, 0.8459]
inp 0.7807 [0.7250, 0.8364]
pin 0.7818 [0.7092, 0.8588]
pni 0.6611 [0.6328, 0.6874]
2u-DNF 0.6539 [0.5996, 0.7104]
up-DNF 0.8364 [0.7683, 0.8980]
Table 5: Accuracy computed between the curated and embedding-based preference sets, broken down by query type.
@@ Appendix C Dataset statistics
We consider 14 types of complex queries over KGs introduced in prior work (Ren and Leskovec, 2020), which include conjunctions, disjunctions, and negation. These queries can be represented using query graphs, which we illustrate in Fig. 6.
@@ Appendix D Experimental details
We implement all of our experiments using PyTorch (Paszke et al., 2019), and run them on a machine with one NVIDIA A100 GPU and 120GB of RAM.
@@
We make use of QTO as the base query answering model Bai et al. (2023). QTO relies on a neural link predictor, for which we use ComplEx (Trouillon et al., 2016; Lacroix et al., 2018) trained with an auxiliary relation prediction objective (Chen et al., 2021). We use an embedding dimension of 1,000, learning rate of 0.1, batch size 1,000, and a regularization weight of 0.05.
@@
We tune the hyperparameters of this method using grid search, where $\alpha$ can take values in $\{0.25,0.5,0.75\}$, and $\beta$ in $\{-0.5,0,0.5\}$. We select the best values based on performance on the validation set. We define performance as the sum of pairwise accuracy and ranking performance.
@@
We describe the architecture of NQR in Table 7. We select the best learning rate in $\{1\times 10^{-5},1\times 10^{-4},1\times 10^{-3}\}$.
Table 7: Layer-wise description of the model architecture in NQR. The embedding dimension is 1,000, and in the input layers we concatenate scalars (the current entity score, and the mean similarity scores with the positive and negative preference sets), indicated as +3+3.
@@
Linear + ReLU 1000 + 3 256
Linear + sigmoid 256 2
@@ Appendix E Additional results
We present here results over all types of complex queries we consider in our experiments. As in our main experiments, we compute the metrics for preference sets of increasing sizes, from 1 up to 10. Here, we report the metrics averaged over all 10 steps. We present results categorized by query structure (illustrated in Fig. 6) in Tables 8 and 9.
@@
E.1 Results on curated queries
We additionally evaluate the models on the subset of queries and manually curated preference sets obtained during the validation step described in Appendix˜B. The performance plots when increasing the preference set size are shown in Figure˜7. Given the smaller sample size in this curated subset, we obtain wider confidence intervals. However, we observe that the general trends observed from our experiments using larger preference sets derived from embeddings hold.
We present more detailed results on the subset of curated queries in Table˜10. When comparing these results with performance on preference sets derived from embeddings (see Table˜8), we observe broadly consistent trends in the relative performance of the methods, suggesting that results on the generated preference sets are reasonably predictive of behavior on the curated ones.
In particular, the Cosine baseline achieves the highest pairwise accuracy on both datasets, with a clear margin over the other methods, indicating that similarity-based reranking aligns well with both synthetic and human preferences. NQR consistently ranks second in pairwise accuracy and exhibits the strongest performance in ranking metrics: it achieves the best average MRR on both datasets and remains competitive in NDCG@10, often close to the Cosine baseline.
Importantly, the relative ordering of methods is largely preserved between datasets: LightGBM provides moderate performance, while Cosine and NQR dominate. At the same time, the curated dataset generally yields higher pairwise accuracy and ranking scores for the top-performing methods, suggesting that certain preference sets derived from embeddings may be harder to model due to imperfect clusters. Overall, these patterns indicate a meaningful correlation between the synthetic and curated evaluations, supporting the use of our setup for generating preference sets as a proxy for estimating real-world performance.
Figure 7: Results on interactive query answering with soft preferences on the subset of manually curated preference sets from the FB15k237 dataset.
Table 10: Per query type results of interactive query answering with feedback on the subset of manually curated preference sets from the FB15k237 dataset.
Method 1p 2p 3p 2i 3i ip pi 2in 3in inp pin pni 2u up avg
Average Pairwise Accuracy
Unconstrained 39.66 54.92 53.16 42.82 55.02 58.34 55.71 58.23 39.35 52.46 40.32 34.87 42.43 44.96 48.02
LightGBM 73.41 73.46 66.20 61.90 73.47 55.11 62.31 83.61 66.28 53.00 63.46 63.01 58.37 66.54 65.72
Cosine 87.68 92.27 95.26 89.09 86.18 91.49 93.35 91.38 82.53 89.34 91.20 89.31 87.62 91.09 89.84
NQR 76.04 90.40 93.59 79.38 76.49 88.58 88.12 90.67 77.19 86.44 85.03 79.88 81.24 84.80 84.13
Average MRR
Unconstrained 19.22 17.46 15.54 30.45 40.27 24.53 20.59 8.04 14.69 6.97 7.28 1.93 14.51 7.36 16.35
LightGBM 16.60 18.76 19.57 27.32 37.69 26.02 14.22 17.35 20.61 10.22 11.52 6.89 14.15 15.01 18.28
Cosine 22.90 18.93 23.00 37.61 38.16 26.90 21.77 19.09 22.03 19.92 17.28 10.49 22.08 22.38 23.04
NQR 27.54 23.24 23.57 37.50 42.10 27.17 25.02 20.26 20.55 20.03 15.89 4.53 20.67 19.37 23.39
Average NDCG@10
Unconstrained 32.99 28.45 27.83 38.17 58.53 37.63 35.48 18.97 27.06 16.39 17.51 4.05 24.15 12.58 27.13
LightGBM 43.87 39.68 40.80 47.51 65.06 43.67 37.18 44.33 46.48 25.32 29.57 24.32 33.94 37.47 39.94
Cosine 60.65 50.59 59.01 63.31 66.74 57.80 56.59 52.00 53.93 55.02 53.09 39.22 56.85 56.53 55.81
NQR 59.27 56.98 58.29 59.97 66.56 57.56 59.64 52.46 49.22 53.78 49.10 17.79 51.53 48.76 52.92